### v1

# MovSemCL: Movement-Semantics Contrastive Learning for Trajectory Similarity

Abstract
Trajectory similarity computation is fundamental functionality that is used for, e.g., clustering, prediction, and anomaly detection. However, existing learning-based methods exhibit three key limitations: (1) insufficient modeling of trajectory semantics and hierarchy, lacking both movement dynamics extraction and multi-scale structural representation; (2) high computational costs due to point-wise encoding; and (3) use of physically implausible augmentations that distort trajectory semantics. To address these issues, we propose MovSemCL, a movement-semantics contrastive learning framework for trajectory similarity computation. MovSemCL first transforms raw GPS trajectories into movement-semantics features and then segments them into patches. Next, MovSemCL employs intra- and inter-patch attentions to encode local as well as global trajectory patterns, enabling efficient hierarchical representation and reducing computational costs. Moreover, MovSemCL includes a curvature-guided augmentation strategy that preserves informative segments (e.g., turns and intersections) and masks redundant ones, generating physically plausible augmented views. Experiments on real-world datasets show that MovSemCL is capable of outperforming state-of-the-art methods, achieving mean ranks close to the ideal value of 1 at similarity search tasks and improvements by up to 20.3% at heuristic approximation, while reducing inference latency by up to 43.4%.
Codehttps://github.com/ryanlaics/MovSemCL
Introduction
The proliferation of GPS-enabled devices enabled massive collections of vehicle trajectory data, enabling applications such as ride-sharing, logistics, and urban analytics (zheng2015trajectory). A fundamental operation underlying these applications is trajectory similarity computation, which quantifies the similarity between trajectories. This functionality is key to enabling a variety of trajectory applications, like similarity search (su2020making), route recommendation (chen2020learning), and mobility prediction (feng2018deepmove).
Using rigid geometric similarity measures (hausdorff1914grundzuge; frechet1906quelques), traditional trajectory similarity computation methods (cormode2007string; keogh2005exact; vlachos2002discovering) are computationally expensive and ignore underlying semantics. Recent learning-based methods embed trajectories into latent vector spaces using neural architectures such as RNNs (liu2016predicting; li2024clear; deng2022efficient), CNNs (yao2018trajectory; chang2024revisiting), and Transformers (xu2020trajectory; chang2023trajcl), enabling efficient similarity computation. However, these methods still face three key limitations:
Figure 1: Pipeline of MovSemCL. The framework addresses three limitations: Movement-Semantics Encoding extracts movement dynamics (L1), Hierarchical Semantics Encoding captures multi-scale patterns with reduced complexity (L1, L2), and Semantics-Aware Contrastive Learning uses curvature-guided augmentation (L3).
Limitation 1 (L1): Insufficient modeling of trajectory semantics and hierarchy. Trajectories contain movement-semantics and are hierarchical in nature—points form maneuvers, maneuvers compose travels—requiring both semantic feature extraction and hierarchical modeling of fine-grained local patterns and coarse-grained global patterns. Existing methods (liu2016predicting; li2024clear; deng2022efficient; chang2023trajcl; xu2020trajectory) treat trajectories as flat sequences of raw coordinates, failing to capture both movement dynamics and multi-scale structure.
Limitation 2 (L2): Computational inefficiency for trajectories with many locations. Real-world trajectories often contain hundreds of points. RNN-based methods (liu2016predicting; li2024clear; deng2022efficient) lack parallelization, while Transformer-based methods (xu2020trajectory; chang2023trajcl) scale quadratically with sequence length, forcing lossy downsampling that compromises movement fidelity.
Limitation 3 (L3): Semantically-unaware augmentations in contrastive learning. Existing contrastive methods (chang2023trajcl; li2022cltsim) apply generic augmentations that create physically impossible trajectories—randomly masking points causes spatial jumps, while uniform sampling destroys critical movement patterns like turns and intersections, corrupting signals that are important for learning.
To address these limitations, we introduce MovSemCL, a movement-semantics contrastive learning framework for trajectory similarity computation. As shown in Figure 1, MovSemCL transforms raw GPS sequences into interpretable movement features (see Section Movement-Semantics Encoding), segments them into semantically coherent patches, and encodes both local and global patterns through dual-level attention (see Section Hierarchical Semantics Encoding). In addition, MovSemCL features a curvature-guided augmentation strategy that generates physically plausible trajectory views by masking redundant segments and preserving behaviorally salient ones (see Section Semantics‑Aware Contrastive Learning).
Our main contributions are as follows:
  • We propose MovSemCL, a movement-semantics contrastive learning framework for trajectory similarity computation that captures movement dynamics.
  • We design three key components: movement-semantics encoding captures rich movement semantics (L1), hierarchical patch-based encoding reduces computational complexity from quadratic to near-linear (L1, L2), and curvature-guided augmentation (CGA) generates physically plausible, semantics-aware trajectory augmentations for robust learning (L3).
  • Extensive experiments on real-world datasets demonstrate that MovSemCL achieves up to 72.6% more accurate similarity search and 43.4% faster inference compared to state-of-the-art baselines.
Problem Formulation
GPS Trajectory
A GPS trajectory $\mathcal{T}$ is a sequence of timestamped locations:
$\mathcal{T}=\langle(s_{0},t_{0}),(s_{1},t_{1}),\cdots,(s_{L-1},t_{L-1})\rangle,$ (1)
where $s_{i}=(\mathrm{lon}_{i},\mathrm{lat}_{i})$ are spatial coordinates, $t_{i}$ is a timestamp, and $L$ is the trajectory length.
Trajectory Similarity Computation
Given a collection $\mathcal{D}=\{\mathcal{T}_{1},\cdots,\mathcal{T}_{N}\}$ of GPS trajectories with varying lengths, the objective is to learn a representation function $f:\mathcal{D}\rightarrow\mathbb{R}^{d}$ that maps a trajectory $\mathcal{T}\in\mathcal{D}$ to a fixed-dimensional embedding vector $\mathbf{z}=f(\mathcal{T})$, facilitating efficient similarity computation. The similarity between trajectories $\mathcal{T}_{i}$ and $\mathcal{T}_{j}$ is then defined as the Cosine similarity of their embeddings:
$\text{sim}(\mathcal{T}_{i},\mathcal{T}_{j})=\frac{\mathbf{z}_{i}\cdot\mathbf{z}_{j}}{\|\mathbf{z}_{i}\|\|\mathbf{z}_{j}\|}$ (2)
Methodology
Overall Architecture
As shown in Figure 1, MovSemCL encompasses three stages: (1) Movement-Semantics Encoding transforms raw GPS data into movement-semantics features; (2) Hierarchical Semantics Encoding segments trajectories into patches for hierarchical modeling via dual-level attention; (3) Semantics-Aware Contrastive Learning leverages physically plausible augmentations and a contrastive loss to focus learning on behaviorally informative segments.
Movement-Semantics Encoding
Figure 2: Overview of Movement-Semantics Encoding.
As illustrated in Figure 2, the proposed Movement-Semantics Encoding facilitates modeling of trajectory semantics by first normalizing GPS coordinates, extracting movement dynamics features (displacement vectors and heading angles), constructing trajectory-induced spatial graphs via Node2Vec (grover2016node2vec) to encode spatial context into spatial cell embeddings, and composing all features into movement-semantics representations (L1).
Coordinate Normalization
Raw GPS coordinates in the WGS84 coordinate system (longitude, latitude) are not suitable for direct distance computation due to the Earth’s curvature. Therefore, we project GPS coordinates to a planar coordinate system using the Mercator projection (snyder1987map):
$(\mathit{mx}_{i},\mathit{my}_{i})=\mathrm{Mercator}(\mathrm{lon}_{i},\mathrm{lat}_{i})$ (3)
The resulting coordinates are further normalized with respect to the map region of interest:
$(\bar{x}_{i},\bar{y}_{i})=\left(\frac{\mathit{mx}_{i}-x_{\min}^{(m)}}{W_{\mathrm{r}}},\frac{\mathit{my}_{i}-y_{\min}^{(m)}}{H_{\mathrm{r}}}\right),$ (4)
where $W_{\mathrm{r}}$ and $H_{\mathrm{r}}$ are the width and height of the region.
Movement Dynamics Features
We then compute the displacement vectors and heading angle for each point:
$dx_{i}=\begin{cases}0&i=0\\ \bar{x}_{i}-\bar{x}_{i-1}&i>0\end{cases},\quad dy_{i}=\begin{cases}0&i=0\\ \bar{y}_{i}-\bar{y}_{i-1}&i>0\end{cases}.$ (5)
Using these displacement vectors, we calculate the heading angle to capture directional changes:
$\theta_{i}=\begin{cases}\frac{\arctan 2(dy_{i},dx_{i})}{\pi}&\text{if }dx_{i}^{2}+dy_{i}^{2}>\epsilon\\ 0&\text{otherwise,}\end{cases}$ (6)
where $\epsilon=10^{-6}$ prevents division by zero. These features capture directional flow and instantaneous changes, enabling distinction between smooth movements and abrupt maneuvers.
Trajectory-Induced Spatial Graph Construction
To provide spatial context beyond coordinates, we partition the map region into a regular grid of $N_{x}\times N_{y}$ cells and assign each trajectory point to a cell according to its normalized coordinates $(\bar{x}_{i},\bar{y}_{i})$:
$\mathrm{cell}_{i}=\left\lfloor\frac{\bar{x}_{i}}{\Delta_{x}}\right\rfloor\times N_{y}+\left\lfloor\frac{\bar{y}_{i}}{\Delta_{y}}\right\rfloor,$ (7)
where $\Delta_{x},\Delta_{y}$ define the grid resolution and $N_{y}$ is the number of cells along the $y$-axis.
We construct a trajectory-induced directed graph $G=(V,E)$ where each cell is modeled as a node $v\in V$. Given the trajectory collection $\mathcal{D}$, we define edges $e_{ij}\in E$ between cells $i$ and $j$ based on consecutive transitions observed in trajectories:
$e_{ij}=\begin{cases}1&\text{if }\exists\mathcal{T}\in\mathcal{D},\exists\text{ timestamp }t:\\ &\quad\text{cell}_{t}=i\text{ and }\text{cell}_{t+1}=j\\ 0&\text{otherwise}\end{cases}$ (8)
The edge weight $w_{ij}$ represents the number of transitions from cell $i$ to cell $j$:
$w_{ij}=\sum_{\mathcal{T}\in\mathcal{D}}\sum_{t=0}^{L-2}\mathbf{1}[\text{cell}_{t}=i\text{ and }\text{cell}_{t+1}=j].$ (9)
This graph encodes mobility patterns where nodes representing frequently connected cells have large weights, reflecting common movements, while pairs of nodes with no edges represent cells with no immediate movement between them. We apply Node2Vec (grover2016node2vec) to learn a structural embedding for each cell:
$\mathbf{ST}_{i}=\mathrm{Node2Vec}(G,\mathrm{cell}_{i})\in\mathbb{R}^{d_{\mathrm{se}}}$ (10)
Feature Composition
We concatenate all features at each point to form the final movement-semantics representation:
$\mathbf{f}_{i}=[dx_{i},dy_{i},\theta_{i},\mathbf{ST}_{i}]\in\mathbb{R}^{d_{\mathrm{in}}},$ (11)
where $d_{\mathrm{in}}=3+d_{\mathrm{se}}$. Here, the first three elements encode local movement, while $\mathbf{ST}_{i}$ captures spatial context.
Hierarchical Semantics Encoding
Figure 3: Comparison of attention mechanisms.
Conventional flat sequence models struggle to capture both local motion details and global intent (L1) and scale poorly to trajectories with many locations (L2). Our hierarchical encoding builds on the movement-semantics features to address L1 and L2, enabling efficient modeling of trajectory semantics while reducing computational complexity.
Patch Construction
Given an enriched feature sequence $\langle\mathbf{f}_{i}\rangle$ from the previous stage, we partition it into $M=\lceil L/P\rceil$ patches, where $P$ is the patch length and $M$ is the number of patches in the trajectory. We obtain the following representation, where a patch $\mathbf{P}_{j}\in\mathbb{R}^{P\times d_{\mathrm{in}}}$ serves as a locally coherent movement unit:
$\mathcal{P}=\langle\mathbf{P}_{1},\mathbf{P}_{2},\cdots,\mathbf{P}_{M}\rangle$ (12)
This patch-based approach reduces computational complexity from $O(L^{2})$ in conventional flat sequence models to $O(M^{2}+P^{2})$, where $M\ll L$ for typical trajectory lengths, effectively addressing the scalability issues caused by trajectories with many locations (L2). Padding and binary masks are used to support variable-length trajectories, following a previous study (chang2023trajcl).
Intra-Patch Attention
Within each patch, we employ a self-attention layer to capture patterns unique to each local segment. The intra-patch attention outputs $\mathbf{H}_{j}^{\mathrm{intra}}\in\mathbb{R}^{P\times d_{h}}$ are computed as follows.
$\mathbf{H}_{j}^{\mathrm{intra}}=\text{SelfAttn}_{\mathrm{intra}}(\mathbf{P}_{j}+\mathbf{PE}_{\mathrm{local}}),$ (13)
where $\mathbf{PE}_{\mathrm{local}}$ provides local positional encoding. To summarize each patch as a fixed-length embedding, we apply masked average pooling over valid (non-padded) positions:
$\mathbf{h}_{j}=\frac{1}{\sum_{k=1}^{P}(1-m_{j,k}^{\mathrm{intra}})}\sum_{k=1}^{P}(1-m_{j,k}^{\mathrm{intra}})\cdot\mathbf{H}_{j,k}^{\mathrm{intra}},$ (14)
where $m_{j,k}^{\mathrm{intra}}$ is a binary mask indicating padding.
Inter-Patch Attention
The patch embedding $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{M}]^{T}\in\mathbb{R}^{M\times d_{h}}$ is then processed by an inter-patch self-attention layer, capturing long-range dependencies and global trajectory intent:
$\mathbf{H}^{\mathrm{inter}}=\text{SelfAttn}_{\mathrm{inter}}(\mathbf{H}+\mathbf{PE}_{\mathrm{global}}),$ (15)
where $\mathbf{PE}_{\mathrm{global}}$ encodes patch-level position information and padded patches are excluded from the computation.
Trajectory Embedding
Finally, we aggregate the outputs of the inter-patch attention to obtain a compact, expressive trajectory embedding:
$\mathbf{z}=\frac{1}{\sum_{j=1}^{M}(1-m_{j}^{\mathrm{inter}})}\sum_{j=1}^{M}(1-m_{j}^{\mathrm{inter}})\cdot\mathbf{H}_{j}^{\mathrm{inter}},$ (16)
where $m_{j}^{\mathrm{inter}}$ masks out padded patches. The resulting embedding $\mathbf{z}\in\mathbb{R}^{d_{h}}$ robustly preserves both local and global movement semantics of a trajectory.
Semantics‑Aware Contrastive Learning
Figure 4: Comparison of trajectory masking strategies.
While the movement-semantics and hierarchical encodings enable rich trajectory modeling, it is crucial that the model learning focuses on behaviorally informative segments—such as sharp turns or rare maneuvers—rather than overfitting to redundant, straight-line fragments (L3).
To achieve this, we propose a semantics-aware contrastive learning framework. For each trajectory, we generate two different augmented views using Curvature-Guided Augmentation (CGA) to create a positive pair, while embeddings of other trajectories in a batch serve as negative samples. This design ensures physically plausible and semantically meaningful augmentations, enhancing the learning by considering behaviorally relevant patterns.
Curvature-Guided Augmentation
Effective contrastive learning hinges on generating semantically consistent and physically plausible augmented views. MovSemCL generates two different augmented views of each trajectory using Curvature-Guided Augmentation (CGA), which form positive pairs for contrastive learning. This augmentation masks trajectory points with probabilities inversely proportional to their local curvature: high-curvature regions (e.g., turns or intersections) are preferentially retained, while straight-line segments are preferentially dropped. Compared to naive random or block masking (Figures 4a–b), CGA produces views that preserve behaviorally informative patterns better, while avoiding spatial discontinuities (Figure 4c). The CGA procedure is detailed in the appendix.
Contrastive Objective
To learn a model, we adopt the MoCo contrastive learning framework (he2020momentum), which maintains a dynamic queue of negative embeddings. Given a query embedding $\mathbf{z}_{q}$, its corresponding positive key $\mathbf{z}_{k}$, a set of negatives $\{\mathbf{z}_{n}^{-}\}_{n=1}^{N}$, and temperature $\tau$, the contrastive loss is formulated as:
$\mathcal{L}=-\frac{\operatorname{sim}\!\bigl(\mathbf{z}_{q},\mathbf{z}_{k}\bigr)}{\tau}+\log\!\Bigl(e^{\operatorname{sim}\!\bigl(\mathbf{z}_{q},\mathbf{z}_{k}\bigr)/\tau}+\sum_{n=1}^{N}e^{\operatorname{sim}\!\bigl(\mathbf{z}_{q},\mathbf{z}_{n}^{-}\bigr)/\tau}\Bigr),$ (17)
where $\operatorname{sim}(\mathbf{u},\mathbf{v})=\mathbf{u}^{\!\top}\mathbf{v}/(\|\mathbf{u}\|\|\mathbf{v}\|)$. Only the query encoder is updated via backpropagation; the key encoder is updated using an exponential moving average for stability.
Experiments
To assess the capabilities of MovSemCL, we design experiments to answer the following research questions:
  • RQ1 (Effectiveness): How does MovSemCL perform at trajectory similarity computation?
  • RQ2 (Robustness): How robust is MovSemCL under real-world conditions with noisy and degraded trajectories?
  • RQ3 (Versatility): Can MovSemCL approximate time-consuming heuristic similarity measures with fine-tuning?
  • RQ4 (Efficiency): What are the computational advantages of MovSemCL for large-scale trajectory processing?
  • RQ5 (Component Analysis): How do the components of MovSemCL contribute to its performance?
  • RQ6 (Hyperparameter Sensitivity): How do key hyperparameters affect MovSemCL and what are the best settings?
Experimental Setup
Datasets
We use two complementary real-world datasets. Porto: dense taxi trajectories from Porto, Portugal (July 2013–June 2014), containing 48 points and being 6.37 km long on average, representing short-distance urban mobility. Germany: sparse long-distance trajectories across Germany (2006–2013), containing 72 points and being 252.49 km long on average, capturing diverse inter-city travel patterns. Following established evaluation protocols (chang2023trajcl; li2018deep; liu2022cstrm), we eliminate trajectories with less than 20 or more than 200 points, and we exclude those outside valid geographic regions. For fair comparison with prior work (chang2023trajcl), we use a subset of 200 thousand trajectories from the Porto dataset and the full Germany dataset. Each dataset is then split into training/validation/test sets at 70%/10%/20%. From the test set, we reserve 10 thousand samples for the heuristic similarity approximation, partitioned using the same ratio.
Implementation Details
Based on preliminary experiments (see Figure 6), we set the patch size $P=4$, the embedding dimension $d=256$, the hidden dimension $d_{h}=256$, and we train for 20 epochs with early stopping. We use the Adam optimizer with a learning rate $1e^{-4}$, batch size 128, and temperature $\tau=0.05$ for contrastive learning. For the spatial discretization, we employ grids with cell sizes adapted to the spatial scale of each dataset: 100 meters for Porto and 1000 meters for Germany. All experiments are conducted on NVIDIA RTX A6000 GPUs.
Trajectory Retrieval Performance (RQ1)
We first investigate MovSemCL’s effectiveness on the core task of trajectory similarity computation.
Experimental Protocol
Following baseline methods (chang2023trajcl), we evaluate trajectory similarity using a query set $\mathcal{Q}$ and database $\mathcal{D}$ created from 100K test trajectories. We randomly sample 1K trajectories and then split each trajectory $\mathcal{T}^{q}$ into odd points $\mathcal{T}_{a}^{q}=[p_{1},p_{3},p_{5},\cdots]$ and even points $\mathcal{T}_{b}^{q}=[p_{2},p_{4},p_{6},\cdots]$. $\mathcal{T}_{a}^{q}$ serves as query and $\mathcal{T}_{b}^{q}$ as ground-truth in database $\mathcal{D}$, which is augmented with random trajectories to form databases of varying sizes. This splitting creates reasonable similar pairs representing the same movement sequence with different sampling offsets. We report the mean rank of the ground-truth $\mathcal{T}_{b}^{q}$ when retrieving similar trajectories for each query $\mathcal{T}_{a}^{q}$, with 1 being the ideal rank and smaller ranks being better.
Results and Analysis
Table 1 presents the mean rank results across different database sizes. MovSemCL consistently achieves the best performance, with mean ranks very close to 1 across all database sizes.
The performance gap reveals fundamental limitations in existing approaches. Traditional geometric methods (EDR, CSTRM, Hausdorff) measure spatial alignment without movement dynamics, causing severe degradation as geometric similarity becomes ambiguous in large databases. RNN-based methods (t2vec, TrjSR, E2DTC) struggle with long-range dependencies and lack parallelization, while the Transformer-based TrajCL treats trajectories as flat sequences, missing hierarchical movement structure. Additionally, TrajCL’s use of random augmentation creates physically implausible trajectories, weakening its contrastive learning. MovSemCL’s near-optimal performance across both dense urban (Porto) and sparse long-distance (Germany) datasets indicates that movement semantics—rather than geometric or temporal patterns alone—provide robust discriminative features that scale effectively. The stability across database sizes reflects a key insight: trajectory similarity is fundamentally about behavioral similarity, which existing methods do not capture comprehensively.
Dataset Method 20K 40K 60K 80K 100K
Porto EDR 8.318 14.398 17.983 22.902 28.753
CSTRM 4.476 7.954 10.630 13.576 16.699
EDwP 3.280 4.579 5.276 6.191 7.346
Hausdorff 3.068 4.014 4.649 5.451 6.376
Fréchet 3.560 4.959 5.968 7.192 8.631
t2vec 1.523 2.051 2.257 2.612 3.068
TrjSR 1.876 2.783 3.208 3.826 4.635
E2DTC 1.560 2.111 2.349 2.731 3.213
CLEAR 1.435 1.593 1.766 1.923 2.077
TrajCL 1.005 1.006 1.006 1.007 1.010
MovSemCL 1.002 1.004 1.005 1.005 1.005
Germany EDR 279.385 558.288 834.208 1108.975 1370.004
CSTRM OOM OOM OOM OOM OOM
EDwP 2.168 2.277 2.371 2.454 2.515
Hausdorff 2.803 3.509 4.206 4.906 5.551
Fréchet 2.581 3.108 3.633 4.113 4.589
t2vec 1.571 1.982 2.387 2.718 3.053
TrjSR 6.517 11.741 16.969 22.182 24.083
E2DTC 3.136 5.156 7.248 9.207 10.956
CLEAR 1.104 1.138 1.177 1.202 1.222
TrajCL 1.012 1.022 1.034 1.040 1.045
MovSemCL 1.002 1.003 1.003 1.005 1.008
Table 1: Mean rank vs. database size (RQ1). Best results are in bold, second-best are underlined. MovSemCL consistently achieves the best mean rank (ideal=1).
Robustness Evaluation (RQ2)
We proceed to examine MovSemCL’s robustness to data degradation that occurs commonly in real-world settings.
Down-sampling Robustness
GPS trajectories often suffer from missing data points due to signal loss, battery constraints, or privacy-preserving sampling. Following prior studies (chang2023trajcl; li2024clear), we down-sample trajectories in $\mathcal{Q}$ and $\mathcal{D}$ by randomly masking points in each trajectory with a probability $\rho_{s}\in[0.1,0.5]$, while keeping $|\mathcal{D}|=100,000$. Table 2 shows that MovSemCL achieves the best performance across all down-sampling rates. Among deep learning methods, TrajCL performs well at low rates but deteriorates significantly as degradation increases (36.352 at 0.5 on Porto). CLEAR, designed with augmentation strategies for robustness, maintains more stable performance on intra-city dataset Porto but degrades on inter-city trajectories like Germany. However, MovSemCL exhibits the best robustness across most settings, clearly demonstrating that movement-semantics modeling provides inherent robustness to data sparsity.
Distortion Robustness
Real GPS data contains coordinate inaccuracies due to sensor noise and atmospheric interference. We follow prior studies (chang2023trajcl; li2024clear) and apply random coordinate shifts to a proportion $\rho_{d}\in[0.1,0.5]$ of trajectory points. Table 3 shows that MovSemCL again exhibits the best robustness, maintaining mean ranks very close to 1 across all distortion rates.
Dataset Method 0.1 0.2 0.3 0.4 0.5
Porto EDR 57.173 203.993 806.033 2286.821 4872.231
CSTRM 24.794 47.137 123.124 257.540 687.262
EDwP 8.442 10.968 18.727 28.394 68.061
Hausdorff 10.026 23.293 56.561 89.827 275.206
Fréchet 10.668 18.516 29.740 93.851 181.271
t2vec 4.786 8.461 19.689 35.219 115.364
TrjSR 7.941 15.746 151.948 549.108 1341.883
E2DTC 5.100 9.385 21.845 39.402 124.320
CLEAR 1.518 1.914 2.792 3.650 6.090
TrajCL 1.026 1.191 1.513 3.847 36.352
MovSemCL 1.018 1.098 1.682 1.961 9.951
Germany EDR 1368.829 1379.489 1375.261 1380.517 1389.433
EDwP 2.173 2.509 2.176 2.191 2.209
Hausdorff 2.514 2.742 4.353 4.448 5.627
Fréchet 2.358 2.492 3.735 3.824 4.642
CSTRM OOM OOM OOM OOM OOM
t2vec 4.453 6.736 9.087 9.470 9.775
TrjSR 24.539 30.318 55.002 68.070 111.175
E2DTC 11.595 13.478 15.843 18.532 19.134
CLEAR 1.265 1.276 1.396 1.460 1.740
TrajCL 1.048 1.050 1.059 1.418 2.045
MovSemCL 1.001 1.008 1.080 1.151 1.265
Table 2: Mean rank vs. down-sampling rate (RQ2).
Dataset Method 0.1 0.2 0.3 0.4 0.5
Porto EDR 28.243 28.498 27.899 28.070 28.932
EDwP 7.591 7.166 7.038 7.235 7.236
Hausdorff 6.549 6.737 6.706 6.592 6.739
Fréchet 8.689 8.854 8.755 8.636 9.083
CSTRM 20.860 20.081 22.081 24.688 26.243
t2vec 3.212 3.487 3.981 3.897 3.999
TrjSR 4.781 5.087 35.144 6.194 7.201
E2DTC 3.348 3.678 4.210 4.129 4.222
CLEAR 1.345 1.313 1.330 1.309 1.356
TrajCL 1.022 1.154 1.076 1.091 1.039
MovSemCL 1.004 1.012 1.005 1.006 1.004
Germany EDR 1373.985 1372.984 1373.981 1373.966 1373.944
EDwP 2.488 2.489 2.492 2.489 2.489
Hausdorff 5.587 5.576 5.573 5.566 5.568
Fréchet 4.631 4.625 4.609 4.625 4.612
CSTRM OOM OOM OOM OOM OOM
t2vec 3.863 3.976 4.903 3.580 3.625
TrjSR 27.146 27.156 27.032 26.935 27.035
E2DTC 10.946 11.161 10.940 11.275 10.693
CLEAR 1.223 1.194 1.209 1.190 1.233
TrajCL 1.049 1.051 1.049 1.062 1.054
MovSemCL 1.008 1.008 1.008 1.008 1.007
Table 3: Mean rank vs. distortion rate (RQ2).
Heuristic Similarity Approximation (RQ3)
Beyond similarity search, we evaluate whether or not MovSemCL’s learned representations can effectively approximate traditional distance measures, thereby assessing the versatility of its movement-semantics embeddings.
Experimental Protocol
Following the prior work (chang2023trajcl), we fine-tune the pre-trained MovSemCL encoder with a two-layer multilayer perceptron head to predict EDR, EDwP, Hausdorff, and Fréchet distances using the MSE loss. This setup tests whether the movement-semantics representations capture sufficient information to approximate these time-consuming heuristic measures. We compare against both fine-tuning models and supervised methods trained specifically for distance prediction, including TrajSimVec (zhang2020trajectory), TrajGAT (yao2022trajgat), and T3S (yang2021t3s). We report HR@$k$ (Hit Ratio at $k$), the proportion of ground-truth top-$k$ similar trajectories correctly identified in predicted top-$k$ results, and R5@20 (Recall of top-5 in top-20), measuring recall of returning ground-truth top-5 trajectories in the top-20 results.
Results
Table 4 shows that MovSemCL achieves an average rank of 1 across all measures. Notably, MovSemCL demonstrates improvements over TrajCL: 20.3% improvement on EDR, 1.7% on Hausdorff, and 1.8% on Fréchet for the HR@5 metrics. The consistently high R5@20 scores (>0.95>0.95) for Hausdorff and Fréchet indicate that the movement-semantics representations does very well at capturing the geometric properties these measures emphasize.
Method EDR EDwP Hausdorff Fréchet Average
rank
HR@5 HR@20 R5@20 HR@5 HR@20 R5@20 HR@5 HR@20 R5@20 HR@5 HR@20 R5@20
t2vec 0.125 0.164 0.286 0.399 0.518 0.751 0.405 0.549 0.770 0.504 0.651 0.883 6
TrjSR 0.137 0.147 0.273 0.271 0.346 0.535 0.541 0.638 0.880 0.271 0.356 0.523 9
E2DTC 0.122 0.157 0.272 0.390 0.514 0.742 0.391 0.537 0.753 0.498 0.648 0.879 7
CSTRM 0.138 0.191 0.321 0.415 0.536 0.753 0.459 0.584 0.813 0.421 0.557 0.768 4
CLEAR 0.158 0.207 0.351 0.487 0.594 0.831 0.596 0.687 0.936 0.583 0.709 0.937 3
TrajCL 0.172 0.222 0.376 0.546 0.646 0.881 0.643 0.721 0.954 0.618 0.740 0.955 2
MovSemCL 0.207 0.308 0.487 0.536 0.642 0.873 0.654 0.742 0.970 0.629 0.741 0.957 1
TrajSimVec 0.119 0.163 0.285 0.172 0.253 0.390 0.339 0.429 0.543 0.529 0.664 0.894 10
TrajGAT 0.090 0.102 0.184 0.201 0.274 0.469 0.686 0.740 0.969 0.362 0.403 0.704 8
T3S 0.140 0.192 0.325 0.377 0.498 0.702 0.329 0.482 0.668 0.595 0.728 0.946 5
Table 4: HR@5, HR@20, and R5@20 of self-supervised and supervised methods to approximate heuristic measures on Porto (RQ3).
Efficiency Analysis (RQ4)
To assess deployment feasibility, we compare MovSemCL’s computational efficiency with that of the SOTA method TrajCL, which also adopts self-attention-based encoder.
Inference Efficiency
Table 5 compares MovSemCL with TrajCL under identical settings (embedding size = 256, batch size = 128, same hardware). MovSemCL achieves substantial improvements across all metrics: 41.2% fewer FLOPs, 43.4% faster inference, and 76.6% higher throughput. These gains arise from replacing the global $O(L^{2})$ attention with block-wise computations of $O(L\cdot P+P^{2})$. The dominant term depends on the trajectory length: when $L<P^{3}$, the $O(L\cdot P)$ term governs, yielding near-linear scaling; otherwise, the bounded $O(P^{2})$ term dominates. In both cases, the effective cost is significantly lower than full $O(L^{2})$ attention.
Figure 5: Scalability analysis (RQ4).
Scalability Analysis
Figure 5 reports MovSemCL’s scalability across varying workloads. The total processing time scales linearly with the number of trajectories, confirming performance without computational bottlenecks. Crucially, the per-sample latency remains remarkably stable (at $\sim$3.4ms) regardless of the dataset size, indicating consistent high efficiency from small batches to large-scale processing.
Metric TrajCL MovSemCL Improv.
FLOPs (M) 158.69 93.34 41.2%
Latency (ms) 6.08 3.44 43.4%
Throughput (samples/s) 164.46 290.41 76.6%
Table 5: Efficiency comparison (RQ4). MovSemCL achieves substantial efficiency improvements over TrajCL.
Ablation Study (RQ5)
To evaluate the contribution of each system component, we remove the movement-semantics encoding (MSE, using only cell embedding), the hierarchical semantics encoding (HSE, using flat sequence processing), and the curvature-guided augmentation (CGA, using random point mask).
Results
Table 6 reveals a clear hierarchy: MSE has the most critical impact, with its removal causing severe degradation, offering evidence that movement semantics are essential. CGA provides moderate but consistent improvements, validating the semantics-aware masking. HSE contributes the least but still provides meaningful improvements, confirming that hierarchical encoding is beneficial.
Method Porto Germany
20K 100K 20K 100K
MovSemCL-MSE 1.521 3.045 1.595 4.122
MovSemCL-HSE 1.005 1.012 1.010 1.039
MovSemCL-CGA 1.033 1.098 1.180 1.234
MovSemCL 1.002 1.005 1.002 1.008
Table 6: Ablation study (RQ5). MSE is the most impactful module, and all modules are effective.
Hyperparameter Study (RQ6)
Finally, we examine MovSemCL’s sensitivity to key hyperparameters on Porto to provide implementation guidance.
Training Epoch
Figure 6(a) shows that MovSemCL converges rapidly within 10 epochs, with stable performance extending to 20 epochs without overfitting. This rapid convergence reduces training costs while ensuring reliability.
Training Trajectory Size
Figure 6(b) shows that performance gains plateau around 20K trajectories for standard conditions, with degraded conditions requiring additional data for optimal performance. This provides practical guidance for minimum training data requirements.
Embedding Dimension
Figure 6(c) reveals optimal performance at dimensionalities in the range 256–512, with substantial improvement from smaller dimensions to 256, followed by stabilization. This indicates that the hierarchical encoding efficiently utilizes the embedding space without excessive parameters.
Patch Size
Figure 6(d) shows that patch size 4 achieves optimal performance. Smaller patches lack sufficient context, while larger patches dilute movement-semantics.
Figure 6: Hyperparameter study (RQ6).
Related Work
Traditional Trajectory Similarity Computation
Traditional methods rely on geometric and statistical principles. Alignment-based approaches adapt string matching algorithms, such as EDR (chen2005robust) which extends edit distance with spatial thresholds. Geometric methods include Hausdorff distance (hausdorff1914grundzuge), measuring maximum point-to-nearest-neighbor distance, Fréchet distance (frechet1906quelques), considering spatio-temporal ordering, and EDwP (ranu2015indexing) projecting trajectories onto a road network. However, these methods are computationally expensive and ignore movement semantics.
Learning-Based Trajectory Representation
Deep learning enables vector representations of trajectories that capture complex spatial-temporal patterns. Early approaches like t2vec (li2018deep) use sequence-to-sequence autoencoders, while TrjSR (cao2021accurate) and E2DTC (fang20212) employ recurrent architectures with attention mechanisms. Recent contrastive learning methods include TrajCL (chang2023trajcl) with trajectory-specific augmentations and dual-feature attention, and CLEAR (li2024clear) with multi-positive contrastive learning. However, existing methods struggle with hierarchical movement modeling, computational efficiency for trajectories with many locations, and semantically-aware augmentations—limitations that MovSemCL addresses.
Conclusion
We present MovSemCL, a movement-semantics contrastive learning framework that enriches GPS trajectories with movement-semantics features and uses hierarchical patch-based encoding with curvature-guided augmentation. Experiments show that MovSemCL is capable of outperforming state-of-the-art methods, reporting mean ranks close to 1 in similarity search, up to 20.3% improvements in heuristic approximation, and up to 43.4% faster inference, while maintaining the best robustness to data degradation.
Acknowledgments
This work was supported by Independent Research Fund Denmark (No. 1032-00481B).